OpenStudy (shelby1290):

Solve by substitution: a+b=5 3a+4b=15

OpenStudy (sleepyjess):

What do you want to solve for first, a or b?

OpenStudy (shelby1290):

I'll go with a

OpenStudy (sleepyjess):

Ok, so we need to get b on one side. The first equation will probably be easiest to do this with.

OpenStudy (shelby1290):

so like: b=5a ?

OpenStudy (sleepyjess):

Close, good try :) It would actually be b = 5 - a since we need to subtract a

OpenStudy (shelby1290):

alright

OpenStudy (sleepyjess):

Now, substitute that in for b in the second equation.

OpenStudy (shelby1290):

3(5-a)+4b=15

OpenStudy (sleepyjess):

That equation equals b, not a :)

OpenStudy (shelby1290):

okay 3(5-b)+4b=15 ...right?

OpenStudy (sleepyjess):

3a + 4(5 - a) = 15

OpenStudy (shelby1290):

ohhh

OpenStudy (shelby1290):

3a+20-4a=15

OpenStudy (sleepyjess):

Do you get why I put it in for b though?

OpenStudy (shelby1290):

yes i do

OpenStudy (sleepyjess):

that is correct, now just solve for a

OpenStudy (shelby1290):

-1a=15-20

OpenStudy (sleepyjess):

Yep, so -a = -5. Now we need to multiply the entire thing by -1 to get positive a

OpenStudy (shelby1290):

a=5

OpenStudy (sleepyjess):

Yes! Now we can substitute in 5 for a in either of the equations and get b

OpenStudy (shelby1290):

I can't tell if it's easier to plug it into the 1st equation or the 2nd

OpenStudy (sleepyjess):

The first probably, all that is in the first is just subtracting from both sides

OpenStudy (shelby1290):

okay so 5+b=5 bring 5 to the other side.... 5-5=0

OpenStudy (shelby1290):

b=0

OpenStudy (shelby1290):

so therefore the poi is (5,0)

OpenStudy (shelby1290):

@sleepyjess thanks for the help!

OpenStudy (sleepyjess):

yep, that is correct, sorry, was afk for a little bit :)